MsBrains (ms-brains):

CAN SOMEONE PLEASE HELP ME WITH THESE 3 REALLY TRICKY QUESTIONS! I WILL MEDAL AMD FAN! (P.S: If I am not online, please try to answer, I will even freaking pay you owl bucks!)

MsBrains (ms-brains):

Information

MsBrains (ms-brains):

Question 1

MsBrains (ms-brains):

Question 2

MsBrains (ms-brains):

Question 3

OpenStudy (studygurl14):

Only one question at once please

MsBrains (ms-brains):

Okay, Let's do number one than. They are all part of the same thing though, that's why I thought they should be together.

OpenStudy (studygurl14):

Okay, so how many of her fans are old enough to drive?

OpenStudy (studygurl14):

hint: add together all the numbers in the bottom two rows

MsBrains (ms-brains):

6+5+5+1+4+2=23

OpenStudy (studygurl14):

yep

OpenStudy (studygurl14):

Now, how many are NOT old enough to drive?

MsBrains (ms-brains):

18+9 =27

OpenStudy (studygurl14):

yep, now add 27 and 23 to find the total # of fans.

MsBrains (ms-brains):

23 +27 = 50

MsBrains (ms-brains):

I don't think you solve the question like that.... 50 doesn't seem like the correct answer...

OpenStudy (studygurl14):

We're not done. Divide the # of fans able to drive by the total

OpenStudy (studygurl14):

convert decimal to percentage

MsBrains (ms-brains):

23/50= 0.46

OpenStudy (abdullahm):

That's correct, and complete for question #1. But one little problem 0.46 is not a percentage. Can you convert 0.46 into a percentage?

OpenStudy (studygurl14):

^

MsBrains (ms-brains):

46% ?

OpenStudy (abdullahm):

Correct :)

OpenStudy (abdullahm):

And for question #2, there are 1000 tickets sold but only 46% people drive. Multiply 46% by 1000 to find out how many car spots will be needed.

MsBrains (ms-brains):

Okay, so just to get in all together, What do I write in the box as an answer? for #1

OpenStudy (abdullahm):

We can assume that all of the people driving will bring their own car, and the rest will be either getting a ride from one of the 46% people or someone else.

OpenStudy (abdullahm):

For #1, you just write whatever we've done so far. So something like this: 6+5+5+1+4+2=23 23 of her fans are old enough to drive. 18+9 =27 27 of her fans are NOT old enough to drive. 23 + 27 = 50 There were a total of 50 fans surveyed. 23/50 = 46% 46% will need a parking spot.

MsBrains (ms-brains):

Okay, and for #2 46% *1000 =460

OpenStudy (abdullahm):

Exactly. So we would need at least 460 parking spaces. The other 54% would either be driving with these 46% people who can drive or arrange a drive another way. :)

OpenStudy (studygurl14):

Looks good. :)

MsBrains (ms-brains):

Alright, so for #2, I would write something as "we would need at least 460 parking spaces. The other 54% would either be driving with these 46% people who can drive or arrange a drive another way." ?

OpenStudy (abdullahm):

yes, but include the calculations 46% * 1000 = 0.46 * 1000 = 460

OpenStudy (studygurl14):

Don't forget to include the math (46% x 1000)

MsBrains (ms-brains):

Alright, Awesome. Now Last one #3

OpenStudy (abdullahm):

mhmmm, think about the last one. Do you think there will be more or less? We can argue for either side, really.

MsBrains (ms-brains):

I think their will be less.. What do you think?

MsBrains (ms-brains):

Guys?

OpenStudy (abdullahm):

Oh sorry, yes, we could write a strong argument that there would be less cars. What's one reason you could list?

MsBrains (ms-brains):

@AbdullahM @StudyGurl14 Sorry I didn't reply, I lost my internet. Um..1 reason...Hmm. I got nothing. How about you guys?

OpenStudy (studygurl14):

The number of cars would probably be less than the estimate. It is impossible for there to be more than the estimate, because it is impossible for someone to drive two cars at once. But some people will come together, and so, logically, there would be less cars.

MsBrains (ms-brains):

Oh, Okay. Can I use that as my part three? :

OpenStudy (studygurl14):

Yeah. But you should restate it in your own words

OpenStudy (abdullahm):

Exactly what StudyGurl said. Also because if a group of a friends who can drive want to go to the concert, they would most likely car pool instead of all going in different cars. But if you wanted to argue for the opposite side, you could write that the survey could have been flawed. The response of 50 people cannot accurately describe the thousands of people who attend the concert.